# A Taxonomy of Multi-Layered Runtime Guardrails for Designing Foundation Model-Based Agents: Swiss Cheese Model for AI Safety by Design

### v2

@@ IV-C2 Targets for Guardrails
  • Models: Models include FMs and narrow models for specific tasks. Guardrails ensures the outputs generated by models are relevant, appropriate and safe. Also, guardrails oversee the utilization of both FMs and narrow models, preventing misuse and ensuring their application under appropriate conditions [1, 20].
  • External Data: Guardrails enforce stringent monitoring and validation of external data sources, particularly in Retrieval augmented generation scenarios [17]. For example, they can prevent the integration of unverified news that could potentially mislead the system’s outputs, ensuring that only reliable and accurate information is used [73].
  • Non-AI Components: Guardrails also apply to non-AI components, e.g. when making calls APIs [1, 71]. For instance, they can monitor API calling to prevent data leaks [74], such as restricting access to sensitive information.
  • Agent-Specific Targets: Guardrails target following key aspects of agents:
@@
    • Reasoning: Enhancing agents’ reasoning capabilities to ensure accurate analysis and sound decision-making [30].
@@
    • Actions: Monitoring agents’ actions to ensure they are safe and effective in achieving desired outcomes [36].
    • Tools: Overseeing the proper use of tools by agents, including implementing access controls, restricting tool capabilities, and detecting potential vulnerabilities [49, 36].
@@
    • Final Results: Final results are the end outputs generated by agents, which are delivered to users or downstream systems. Guardrails ensure that the final results meet user expectations and comply with ethical guidelines and legal regulations.
IV-C3 Guardrails Scopes
The scope of guardrails in FM-based agents ranges from individual preferences to industry standards, reflecting the multi-layered approach of the Swiss Cheese Model. At the user level, guardrails reflect individual preferences and requirements. This involves adjusting the system’s behavior based on user-defined settings to align outputs with both user expectations and ethical considerations. Incorporating user preferences into guardrails provides a personalized experience while maintaining safety and compliance [55, 69]. Such guardrails ensure that the agents respects user autonomy and produces outputs that are relevant and acceptable.
@@ IV-C3 Guardrails Scopes
Industry-level regulations and standards provide the broader regulatory framework within which FM-based agents must operate. Guardrails designed to comply with these regulations guarantee that the system adheres to industry best practices and legal requirements [16]. Guardrails facilitate simpler auditing and certification processes, ensuring the system remains compliant with evolving regulatory landscapes.
System-level constraints focus on the technical and operational limitations within which the FM-based system operates. Guardrails at this level ensure that the system functions efficiently within these constraints, such as computational and memory limits, while maintaining robustness and reliability [26]. They also ensure that the system’s operations do not exceed predefined thresholds that could lead to performance degradation or security vulnerabilities.
IV-C4 Rules
Guardrails rules can be configured in different ways: including uniform rules, priority-enabled rules, context-dependent rules, and negotiable rules, forming layers of defense as per the Swiss Cheese Model. A uniform strategy applies the same set of guardrails consistently across all scenarios, ensuring simplicity and uniformity [67]. It is particularly effective in environments with stable and well-understood risks. It largely reduces the complexity of managing diverse guardrails [55]. A priority-enabled strategy prioritizes certain guardrails based on the criticality and sensitivity of operations or data. Context-dependent strategies adjust the implementation of guardrails based on the system’s specific operational context. This allows for dynamic adjustments to guardrails in response to changing conditions, user needs, and operational environments [49]. The negotiability of guardrails, categorized into hard and soft, defines the level of flexibility in enforcing rules. Soft guardrails allow adjustments based on context and situational demands, providing a balance between protection and operational flexibility [49]. In contrast, hard guardrails are rigid and non-negotiable, ensuring adherence to critical legal, ethical, or safety standards [12, 32].
IV-C5 Autonomy
Guardrails vary based on the level of autonomy in the system. They can be either automatic or require human intervention. Automatic guardrails function without human oversight, relying on pre-defined rules and real-time monitoring to enforce protections [53]. This is useful in scenarios requiring immediate response to potential threats.
In contrast, human intervention guardrails involve manual oversight, where human operators review and approve critical outputs or actions flagged by the system [32]. This approach is essential in high-stakes environments, such as financial systems, where human judgment can provide an additional layer of scrutiny and ensure compliance with regulatory standards.
IV-C6 Modality
The modality of guardrails refers to the types of data and interactions they manage. Guardrails can be designed for single modal or multimodal systems. Single modal systems operate with one type of data input or output, such as text, image, or audio. For instance, in text-based agent systems, guardrails focus on addressing issues like offensive language, misinformation, and data privacy [49]. In image-based agent systems, they may involve techniques for detecting explicit content or ensuring image quality standards [26].
@@ IV-C6 Modality
IV-C7 Underlying Technique
The underlying techniques of guardrails include rule-based, hybrid, and machine learning models, with each representing a distinct design option to meet specific requirements. These techniques offer the necessary flexibility, adaptability, and robustness to protect the systems [3, 27]
@@ IV-C7 Underlying Technique
Rule-based models utilize predefined rules to monitor and control FM-based system behavior. These models implement strict and deterministic guidelines that the system must follow to ensure compliance with regulatory requirements for data access and processing [49]. They are particularly effective in environments where operational parameters are well-defined and stable. Rule-based agent systems can be updated and are somewhat flexible. However, they may still struggle with unexpected scenarios, such as detecting novel AI-generated content that falls outside predefined rules. This reliance on static rules can limit their adaptability and regular updates are needed [71, 16].
@@
In contrast, machine learning models dynamically adapt and improve guardrails based on new data and scenarios. These models can also learn from historical data and identify patterns that indicate potential risks or compliance issues [64]. Machine learning models can be further subclassed into narrow models and foundation models. Narrow models are specialized systems designed for specific tasks or domains. They require targeted guardrails to address domain-specific risks and compliance needs [15]. Foundation models are large, general-purpose models that serve as the backbone for multiple applications and tasks. These models necessitate comprehensive and scalable guardrails to handle a wide range of risks and compliance issues across different applications [26]. Nevertheless, they can be computationally intensive and require substantial data for training.
@@
Hybrid models integrate rule-based agent systems with the adaptability of machine learning models to respond to new threats and evolving data patterns [53]. For instance, Khorramrouz et al.[59] demonstrate the use of the PaLM 2 framework to process user input and dynamically implement rule-based decisions. This framework tests the system’s limits by iteratively generating toxic content to evaluate PaLM 2’s safety guardrails. However, integrating hybrid models can increase system complexity and create additional challenges [53].
V Discussion
In this section, we discuss the implications of our proposed taxonomy, its alignment with the Swiss Cheese Model for AI safety, and how it addresses the challenges in designing multi-layered runtime guardrails for FM-based agents. We also reflect on the potential impact of our findings on future research and practice in AI safety by design.
V-A Adapting the Swiss Cheese Model for Multi-Layered Guardrails
The Swiss Cheese Model, originally developed in risk management and safety science, illustrates how multiple layers of defense can prevent hazards from leading to adverse outcomes [75]. Each layer has potential weaknesses (holes), but when combined, they provide a robust barrier against failures. In the context of FM-based agents, the Swiss Cheese Model underscores the necessity of implementing multi-layered guardrails to address the unique challenges introduced by these agents’ autonomous and non-deterministic behaviors.
FM-based agents interact with various components during runtime, including goals, prompts, plans, tools, knowledge bases, and intermediate and final results. Each of these components represents a potential point of failure or risk. By adapting the Swiss Cheese Model, we propose that guardrails should be implemented at multiple layers of the agent architecture to effectively mitigate risks. This multi-layered approach ensures that if one layer fails to detect or prevent an undesirable behavior, subsequent layers can compensate, thereby enhancing the overall safety and reliability of the agent.
Implementing guardrails at different levels—such as input processing, model reasoning, action execution, and output generation—aligns with the Swiss Cheese Model’s principle of layered defenses, including overlapping layers. This approach addresses the complex interplay between different components in FM-based agents and provides a systematic framework for AI safety by design. It also highlights the importance of considering both external and internal quality attributes in guardrail design, as well as the trade-offs involved in different design options.
V-B Implications for Practice and Research
Our taxonomy offers practical guidance for developers and researchers in designing multi-layered runtime guardrails for FM-based agents. By categorizing the motivations, quality attributes, and design options, we provide a structured approach to address AI safety concerns effectively. Practitioners can use this taxonomy to make informed architectural decisions, balancing trade-offs between different quality attributes such as accuracy, performance, adaptability, interpretability and so on.
For researchers, our work assists in exploring new techniques and methodologies in guardrail design and implementation. The alignment with the Swiss Cheese Model encourages a holistic view of safety mechanisms, promoting research into how different layers can interact synergistically to enhance agent safety. Furthermore, understanding the relationships between guardrails, the risks they mitigate, and the quality attributes they influence can inform the development of standards and best practices in the field of AI safety.
V-C Threats to Validity
@@
Our study is subject to standard literature search and selection bias threats. We addressed these threats by searching the most commonly used databases in the IT and software engineering domains. We adjusted our search strings several times during the automatic search to maximize the number of relevant articles matching two key concepts: ‘guardrails’ and ‘FM-Based agents’. We also kept our search string generic to search through the titles, abstracts, keywords, and full text of articles to cover the maximum number of relevant papers. We then conducted a manual search on Google Scholar to complement the automatic search using a snowballing strategy. Furthermore, predefined review protocols with detailed inclusion and exclusion criteria helped us reduce bias in selecting primary studies. We applied several quality assessment criteria (shown in section III-F) to estimate the quality of the selected primary studies. Even though the proposed criteria were not too strict, applying them indeed caused several initially selected papers to be excluded. To mitigate the risk of missing important data from the primary studies, we reinstated the excluded papers closely related to the primary studies.

### v3

@@ IV-B2 Targets
  • Final Results: Final results are the end outputs generated by agents, which are delivered to users or downstream systems. Guardrails ensure that the final results meet user expectations and comply with regulations and standards.
@@
  • Reasoning: Checking whether the reasoning is sound [30].
@@
  • Workflows: Managing the exceptions happened during runtime workflow execution [80].
  • Tools: Overseeing the proper use of tools by agents, including implementing access controls, restricting tool capabilities, and detect potential vulnerabilities [49, 36].
  • Knowledge Bases: Guardrails enforce stringent monitoring and validation of external knowledge bases, particularly in retrieval augmented generation scenarios [17]. For example, they can prevent the retrieval of sensitive business data [73].
@@
  • FMs: Guardrails ensures the outputs generated by FMs are relevant, appropriate and safe. Also, guardrails oversee the utilization of FMs, preventing misuse and ensuring their application under appropriate conditions [1, 20].
IV-B3 Rules
Guardrails rules can be configured in different ways: including uniform rules, priority-enabled rules, context-dependent rules, and negotiable rules. A uniform strategy applies the same set of guardrails consistently across all scenarios, ensuring simplicity and uniformity [67]. It is particularly effective in environments with stable and well-understood risks. It largely reduces the complexity of managing diverse guardrails [55]. A priority-enabled strategy prioritizes certain guardrails based on the criticality and sensitivity of operations or data. Context-dependent strategies adjust the implementation of guardrails based on the system’s specific operational context. This allows for dynamic adjustments to guardrails in response to changing conditions, user needs, and operational environments [49]. The negotiability of guardrails, categorized into hard and soft, defines the level of flexibility in enforcing rules. Soft guardrails allow adjustments based on context and situational demands, providing a balance between protection and operational flexibility [49]. In contrast, hard guardrails are rigid and non-negotiable, ensuring adherence to critical legal, ethical, or safety standards [12, 32].
IV-B4 Sources
The source of guardrails in FM-based agents ranges from industry regulations and standards to individual preferences. Industry-level regulations and standards provide the broader regulatory framework within which FM-based agents must operate. Guardrails designed to comply with these regulations guarantee that the system adheres to industry best practices and legal requirements [16]. They facilitate simpler auditing and certification processes, ensuring the agent remains compliant with evolving regulatory landscapes.
@@ IV-B4 Sources
Team-level constraints focus on the technical and operational limitations defined by the development team. Guardrails at this level ensure that the agent functions efficiently within these constraints, such as computational and memory limits, while maintaining robustness and reliability [26]. They also ensure that the agent’s operations do not exceed predefined thresholds that could lead to performance degradation or security vulnerabilities.
From the user perspective, guardrails can reflect individual preferences and requirements. This involves adjusting the agent’s behavior based on user-defined settings to align outputs with both user expectations and ethical considerations. Incorporating user preferences into guardrails provides a personalized experience while maintaining safety and compliance [55, 69]. Such guardrails ensure that the system respects user autonomy and produces outputs that are relevant and acceptable.
IV-B5 Modality
The modality of guardrails refers to the types of data and interactions they manage. Guardrails can be designed for single modal or multimodal systems. Single modal systems operate with one type of data input or output, such as text, image, or audio. For instance, in text-based agents, guardrails focus on addressing issues like offensive language, misinformation, and data privacy [49]. In image-based agents, they may involve techniques for detecting explicit content or ensuring image quality standards [26].
@@ IV-B5 Modality
IV-B6 Underlying models
@@
The underlying techniques of guardrails include rule-based, hybrid, and machine learning models, with each representing a distinct design option to meet specific requirements [3, 27]. Rule-based models utilize predefined rules to monitor and control FM-based agents behavior. These models implement strict and deterministic guidelines that the agent must follow to ensure compliance with regulatory requirements for data access and processing [49]. They are particularly effective in environments where operational parameters are well-defined and stable. Rule-based models can be updated and are somewhat flexible. However, they may still struggle with unexpected scenarios, such as detecting novel AI-generated content that falls outside predefined rules. This reliance on static rules can limit their adaptability, and regular updates are needed [71, 16].
@@
In contrast, machine learning models dynamically adapt and improve guardrails based on new data and scenarios. These models can also learn from historical data and identify patterns that indicate potential risks or compliance issues [64]. Machine learning models can be further classified into narrow models and FMs. Narrow models are specialized systems designed for specific tasks or domains. They require targeted guardrails to address domain-specific risks and compliance needs [15]. FMs are large, general-purpose models that serve as the backbone for multiple applications and tasks. These models necessitate comprehensive and scalable guardrails to handle a wide range of risks and compliance issues across different applications [26]. Nevertheless, they can be computationally intensive and require substantial data for training.
Figure 4: Reference architecture for multi-layered guardrails of FM-based agents.
@@
Hybrid models integrate rule-based approaches with the adaptability of machine learning models to respond to new threats and evolving data patterns [53]. For instance, Khorramrouz et al.[59] demonstrate the use of the PaLM 2 framework to process user input and dynamically implement rule-based decisions. This framework tests the system’s limits by iteratively generating toxic content to evaluate PaLM 2’s safety guardrails. However, integrating hybrid models can increase system complexity and create additional challenges [53].
V Reference Architecture for Designing Multi-layered Runtime Guardrails of Agents
Figure 4 shows the proposed reference architecture for multi-layered runtime guardrails of FM-based agents, which consists of four key parts: (i) external environment, (ii) agent components, (iii) built-in multi-layered runtime guardrails, and (iv) AgentOps infrastructure.
V-A External Environment:
The external environment refers to all entities interacting with the agent, including users, other agents, external tools, and knowledge bases. Users provide goals and contextual inputs that shape the agent’s objectives. To achieve user goals, the agent may utilize context detected in the external environment and interact with other agents, specialized tools, and extensive knowledge bases to perform complex tasks.
V-B Agent Components
Within the agent, there are four primary components: the context engine, reasoning and planning, workflow execution, and memory.
  • Context Engine: The context engine processes multimodal context data from the external environment to enrich the user prompt, helping FMs better understand user goals. A prompt may contain elements such as goals and context. Instead of waiting for users’ instructions, the agent can also proactively make suggestions based on the context it detects, such as screen recordings, mouse clicks, eye tracking data, gestures, and document annotations [3].
  • Reasoning and Planning: After receiving optimized prompts, the reasoning and planning component processes the prompt to determine the most effective way of achieving the specified goal. This process may involve adopting reasoning patterns, such as the chain-of-thought pattern [81], which structures the agent’s thinking into sequential, logical steps that align with the agent’s objectives. A detailed plan is then formulated to outline each step required to accomplish the goal. This includes selecting the appropriate tools, knowledge bases, and agents to carry out each action. The memory component may be integrated to allow the agent to recall previously gathered experience and knowledge to refine the plan.
  • Workflow Execution: The workflow execution component is responsible for executing the sequence of actions outlined by the reasoning and planning component. This component directly interacts with external tools, knowledge bases, and other agents to complete tasks and generate outputs aligned with the user’s goals. The results are returned to the external environment and stored in the agent’s memory for future reference.
  • Memory: The memory component in this architecture stores relevant information from prior interactions, plans, and results. This accumulated knowledge supports continuous learning, enabling the agent to refine its strategies and improve capabilities and skills over time, thereby improving accuracy and minimizing repeated errors.
V-C Multi-layered Runtime Guardrails
Building on the Swiss Cheese Model, we design multi-layered runtime guardrails for FM-based agents, structured around the dimensions of quality attributes, pipelines, and artifacts specified in the taxonomy. In this architecture, each ‘cheese slice’ represents a protective layer within the agent system, addressing quality attributes, pipeline stages, and/or specific artifacts, such as a layer about privacy guardrails for prompts or security guardrails for tools. While each layer contains holes (i.e., potential gaps or weaknesses), where risks might slip through, the holes are positioned differently across layers. Gaps in one layer are often covered by another; thus, even if one layer fails, another can catch and mitigate the issue.
From the perspective of quality attributes (discussed in section IV-A), guardrails can be designed to ensure accuracy, efficiency, privacy, security, safety, fairness, compliance. From the pipelines perspective, guardrails can be applied at multiple stages: the user prompts, intermediate results during workflow executions, and final results generated by the agent.
  • Guardrails for prompts: Analyse incoming user prompts to detect and manage sensitive information, harmful content, misinformation, disinformation, discriminatory language, ensuring the prompt aligns with safety and ethical standards [82].
  • Guardrails for intermediate results: Apply at each step of the workflow to verify that intermediate results are accurate, safe, and responsible, safeguarding the integrity of the process before the final results are produced.
  • Guardrails for final results: Check that the agent’s final outputs are align with the user goals and governance requirements, such as AI safety standard requirements.
Moreover, from the artifacts perspective, guardrails can be enforced on each agent artifact including goals, context, memory, reasoning, plans, tools, knowledge bases, other agents, and FMs. These guardrails ensure that each artifact is within safe and responsible boundaries.
  • Guardrails for goals: Ensure that the goals are achievable, within the agent’s scope, and aligned with governance requirements, including regulatory standards and organizational policies, avoiding goals that may lead to harmful outcomes and potential misuse [85].
  • Guardrails for context: Validate contextual information to ensure it is relevant, accurate, and free from sensitive or misleading information.
  • Guardrails for memory: Ensure that stored past experience is relevant, accurate, and free from any malicious or misleading content, preventing memory poisoning [83] and retaining only useful data for future interactions.
  • Guardrails for reasoning: Check the agent’s reasoning processes to prevent logical errors and ensure the reasoning steps are safe, responsible, and aligned with the user intent.
  • Guardrails for plans: Assess the feasibility, safety, and compliance of the plans generated by the agent, ensuring that each step in the workflow is responsible and does not introduce unnecessary risk. The plan can be made by external verifiers, i.e., external planning tools [79].
  • Guardrails for workflows: Handle the exceptions that arise during the workflow executions by implementing mechanisms like force-failing a step or retrying a tool call [80]
  • Guardrails for external tools: Analyse the quality (e.g. vulnerability [86]) of the external tools to ensure that only approved and safe tools are invoked by the agent.
  • Guardrails for knowledge bases: Verify that the information retrieved from knowledge bases is relevant and ethical (e.g., without any PII information).
  • Guardrails for other agents: Ensure the selected agents have a reliable and safe operational history.
  • Guardrails for FMs: Enforce boundaries on the FM’s non-deterministic outputs, applying modifications or flags as needed.
V-D AgentOps
AgentOps provides a comprehensive infrastructure designed to enable observability [84] for FM-based agents by continuously monitoring and recording runtime data. This infrastructure captures a wide range of data elements, from pipeline execution details and agent artifacts to the specific guardrails applied to the pipeline and artifacts. All these data need to be kept as evidence with metadata such as FM version and the timestamp. The data collected by the AgentOps infrastructure can also feed into multi-layered guardrails to activate the relevant guardrails as needed.
VI Threats to Validity
@@
Our study is subject to standard literature search and selection bias threats. We addressed these threats by searching the most commonly used databases in the IT and software engineering domains. We revised our search strings several times during the automatic search to maximize the number of relevant articles matching two key concepts: ‘guardrails’ and ‘FM-based agents’. We also kept our search string generic to search through the titles, abstracts, keywords, and full text of articles to cover the maximum number of relevant papers. We then conducted a manual search on Google Scholar to complement the automatic search using a snowballing strategy. Furthermore, predefined review protocols with detailed inclusion and exclusion criteria helped us reduce bias in selecting primary studies. We applied several quality assessment criteria to estimate the quality of the selected primary studies. Even though the proposed criteria were not too strict, applying them led to several initially selected papers being excluded. To mitigate the risk of missing important data from the primary studies, we reinstated the excluded papers that were closely related to the primary studies.